An electrical company claims that the average life of the bulbs it manufactures is 1 100 hours with a standard deviation of 200 hours. If a random sample of 110 bulbs is chosen, what is the probability that the sample mean will be between 1120 hours and 1 150 hours?

1
Expert's answer
2022-05-17T11:17:18-0400

Let "X=" the sample mean: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=1100h, \sigma=200h, n=110."


"P(1120<X<1150)=P(Z<\dfrac{1150-1100}{200/\sqrt{110}})"

"-P(Z\le\dfrac{1120-1100}{200/\sqrt{110}})"

"\approx P(Z<2.6220)-P(Z\le1.0488)"

"\approx0.99563-0.85286=0.1428"